Find the inverse of y = (3x) / (x-2) and state the domain of the inverse.
I have tried switching the x's for y. I get stuck at this step of the problem:
{(xy)-(2x)} / 3
I've looked online for inverse help, but nothing makes sense! Everyone just says switch the x for y.
Thanks

Found 2 solutions by RAY100, scott8148:
Answer by RAY100(1637)

starting with ,,,,,,y= (3x)/(x-2)
.
Inverse is,,,,,,x=(3y)/(y-2),,,,,,,just switch x for y and vice verse
.
BUT, now lets put in terms of y=?
.
cross multiply ,,,,x(y-2) = 3y
.
xy -2x =3y
.
xy -3y =2x
.
y(x-3) = 2x
.
y= 2x/(x-3),,,,,,,DOMAIN is all REAL numbers except x=3
.
.
inverses should reflect over the line y=x,,,(45 degree line)
.

Answer by scott8148(5880)

switching x and y ___ x = (3y) / (y - 2)

dividing by x and multiplying by (y - 2) ___ y - 2 = 3y / x

dividing by y ___ 1 - (2 / y) = 3 / x

adding 1 (or x/x) ___ -2 / y = (3 + x) / x

take inverse and multiply by -2 ___ y = -2x / (3 + x)

the domain is all real numbers EXCEPT -3